OpenStudy (mudkillz):

Can I have some help with factoring GCF's for polynomials? Question will be below.

OpenStudy (anonymous):

hey u can use a site called mathway.com to get all ur math done

OpenStudy (mudkillz):

\[(4xy^3 + 8x^2y^4 + 12xy^2)\]

OpenStudy (anonymous):

$\left(4xy\cdot\:3+8x\cdot\:2y\cdot\:4+12xy\cdot\:2\right)$(4xy· 3+8x· 2y· 4+12xy· 2) $\mathrm{Multiply\:the\:numbers:}\:4\cdot\:3=12$Multiply the numbers: 4· 3=12 $=12xy+2\cdot\:4\cdot\:8xy+2\cdot\:12xy$=12xy+2· 4· 8xy+2· 12xy $\mathrm{Multiply\:the\:numbers:}\:8\cdot\:2\cdot\:4=64$Multiply the numbers: 8· 2· 4=64 $=12xy+64xy+2\cdot\:12xy$=12xy+64xy+2· 12xy $\mathrm{Multiply\:the\:numbers:}\:12\cdot\:2=24$Multiply the numbers: 12· 2=24 $=12xy+64xy+24xy$=12xy+64xy+24xy $\mathrm{Add\:similar\:elements:}\:12xy+64xy+24xy=100xy$Add similar elements: 12xy+64xy+24xy=100xy $=100xy$=100xy

OpenStudy (kayders1997):

Whats the biggest number they all have in common

OpenStudy (mudkillz):

I know the gcf, give me a sec to right it. I need one more equivalent equation.

OpenStudy (kayders1997):

Okay

OpenStudy (mudkillz):

\[4xy^2(y^1 + 2x^1y^2 + 3)\]

OpenStudy (kayders1997):

Yes thats right

OpenStudy (mudkillz):

I'm trying to figure out one equivalent form to the gcf I just posted.

OpenStudy (ihelpyoulearn):

If you need more help I can help you :) I've seen a lot of people use this site, I'll post below, that shows you every step too solve these questions. Or if anyone needs step by step help with this! http://www.webmath.com/factorgcf.html